OpenStudy (anonymous):

What is the center and radius of the circle with equation x^2 + y^2 + 18x – 12y – 27 = 0?

OpenStudy (anonymous):

Center: (9, –6); Radius = 144 Center: (–9, 6); Radius = 144 Center: (9, –6); Radius = 12 Center: (–9, 6); Radius = 12

OpenStudy (mathmale):

Hello, NK! Please re-arrange x^2 + y^2 + 18x – 12y – 27 = 0 with all the x-terms first and all the y-terms second, and the constant (now +27) on the right side of your equation. Next move: complete the square. How would you complete the square for x^2 + 18x?

OpenStudy (anonymous):

Would x^2 + 18x + y^2 -12y = 27 be my equation?

OpenStudy (anonymous):

Actually, I'm not quite sure how I'd complete the square.

OpenStudy (mathmale):

Great! You're good at following directions. Now, complete the square of x^2 + 18x. Have you any reference work available (textbook, online learning materials, etc.)?

OpenStudy (mathmale):

I'm sure you're familiar with the general formula for a quadratic function: ax^2 + bx + c = 0. Compare that to your 1x^2 + 18x Comparing coefficients, you see that a=1 and b=18, right?

OpenStudy (anonymous):

Yes

OpenStudy (mathmale):

1. Multiply your "b" by 1/2. 2. square the result. 3. Add the result to your x^2 + 18 x, and then subtract that same result. What do you get?

OpenStudy (anonymous):

x^2 + 18x + 81 = 0 x^2 + 18x = -81 ?

OpenStudy (mathmale):

Make it x^2 + 18x + 81 -18. Just hold that for now. Next, look at the y terms: y^2 -12y. What is "b" in this case?

OpenStudy (anonymous):

-12

OpenStudy (mathmale):

Cool. and what is the SQUARE of "HALF OF b"?

OpenStudy (anonymous):

36?

OpenStudy (mathmale):

this is exactly what you did before.

OpenStudy (mathmale):

Right. So, you take your y^2 - 12y and add 36, then subtract 36. Now you'll have x^2 + 18x + 81 - 81 + y^2 and... what else? Please finish typing your whole equation, including that +27 on the right side.

OpenStudy (anonymous):

x^2 + 18x + 81 - 81 + y^2 -12y + 36 - 36 = 27 Would I subtract the 81 & 36 now?

OpenStudy (mathmale):

Add 81 to both sides of the equation, and then add 36 to both sides of the equation. this will leave you with x^2 + 18x + 81 on the left side, plus your y terms. Agreed? what does your equation look like now?

OpenStudy (anonymous):

x^2 + 18x + 81 + y^2 - 12y + 36 = 144

OpenStudy (whpalmer4):

Yes. Can you group those as (factored) perfect squares now?

OpenStudy (mathmale):

Bill: OpenStudy was down, so nk7 and I were continuing work on this problem thru private messaging.

OpenStudy (mathmale):

@nk7 has already typed this in: "A circle, and (-9, 6). And it would represent the radius I believe. Sorry about missing the "^2" part".

OpenStudy (whpalmer4):

ah, sounds good. I didn't see you here, and figured he shouldn't be kept waiting unnecessarily.

OpenStudy (mathmale):

Many thanks, Bill! @nk7: would you mind summarizing your results? "the given equation represents a circle with center at ( , ) and with radius ( )"

OpenStudy (anonymous):

the given equation represents a circle with center at (-9 , 6 ) and with radius of 12

OpenStudy (mathmale):

Nice work!! Any questions about the "completing the square" process?

OpenStudy (anonymous):

None at all, you explained it quite well. I'm a she by the way :) Thank you both immensely. mathmale, thank you for actually explaining this to me rather than just giving me the answer. I actually understand this concept now! And whpalmer4, thank you for stepping in when you thought no one was around. This helped alot!

OpenStudy (mathmale):

congratulations on having learned so quickly and done so well here. I also want to thank you for expressing your appreciation. Hope we can work together again in the future!